OpenStudy (justin528p):

i dont get these at all Dx pls hurry and reply Indicate in standard form the equation of the line passing through the given point and having the given slope for all. 1.A(5, 5), m = 3 2.D(5, -2) m =2/5 3.G(4, 6), H(1, 5) 4.P(6, 2), Q(8, -4) 5.M(0, 6), N(6, 0) 6.B(6, 2), m =-1/2 7.R(3, 3), S(-6, -6) 8.L(5, 0), M(0, 5) 9.X(0, 6), Y(5, 6) 10.E(3, 5), m = undefined 11.A(4, 1), B(5, 2) 12.E(-2, 2), F(5, 1) 13.C(0, 4), m = 0 14.B(0, 0), C(4, -4) please hurry im at the edge of finishing my class this could determine if i fail after this is a quiz then test i need extra score please

OpenStudy (alekos):

well you better get cracking then!

OpenStudy (anonymous):

I'll do 3 of them for you (there's 3 types of questions here), then you can do the rest. I'll happily check your other answers, but I'm not going to do your test for you. 1. A (5,5), m=3. The equation of a straight line is y=mx+c. We have m and a coordinate (x,y) so we can find c: 5=3*5+c so c=5-15=-10. Answer: y=3x-10. 3. G (4, 6), H (1, 5) We aren't given m so we need to find it using coordinates (x0, y0) (x1, y1) (G and H):\[m=\dfrac{y_1-y_0}{x_1-x_9}=\dfrac{5-6}{1-4}=\dfrac{-1}{-3}=3\]Then like in 1, we use y=mx+c with one of the points, say G, 6=3*4+c so c=6-12=-6. Answer: y=3x-6. 10. E(3, 5), m = undefined Not enough information to find it. y=mx+c for some constant c.

OpenStudy (justin528p):

okay youl check behind me??

OpenStudy (justin528p):

? hello

OpenStudy (anonymous):

Post your answers and I'll check them.

OpenStudy (justin528p):

i dont have them yet i was hoping you would work me through it with me

OpenStudy (anonymous):

I've shown you how to do them. Now repeat what I did for the others. As I said before, I'm not doing this for you, it's a lot of work.

OpenStudy (justin528p):

no im not saying that im saying i do the problem and you tell me what i do right and wrong

OpenStudy (anonymous):

Sure.

OpenStudy (justin528p):

k i think i got this so i take the 5 and -2 and the 1/2 then i make it 5=1/2*-2?

OpenStudy (anonymous):

Yes, but we need to find c. So we have 5=(1/2)*2+c. Rearrange to find C.

OpenStudy (justin528p):

make the c by its self?

OpenStudy (anonymous):

Yep :)

OpenStudy (justin528p):

so-c=(1/2)*2-5

OpenStudy (justin528p):

then -c=1-5

OpenStudy (justin528p):

c=-1+5 c=4?

OpenStudy (anonymous):

Sorry I've just seen you have your gradient wrong and I also wrote 2 instead of -2 before. m =2/5. So \[y=\dfrac{2}{5}x+c\]Using the coordinate (5,-2) we get \[5=\dfrac{2}{5}\times -2+c\]What is C?

OpenStudy (justin528p):

no problem 22.5 is c

OpenStudy (justin528p):

49 seconds ill rebump in

OpenStudy (anonymous):

Not quite.\[c=5-\dfrac{-4}{5}=5+\dfrac{4}{5}=\dfrac{29}{5}\]

OpenStudy (justin528p):

okay can u tell me how that happened lol lots of fractions just hit me

OpenStudy (anonymous):

From before, \[5=\dfrac{2}{5}\times -2+c\]\[5=\dfrac{-2\times 2}{5}+c\]\[5=\dfrac{-4}{5}+c\]\[c=5-\dfrac{-4}{5}\] then just look at my last post.

OpenStudy (justin528p):

oh okay i get it i think i can tackle the rest of these alone thank you im a fan!!

OpenStudy (anonymous):

Glad to hear it, let me know if you have any more problems and I'll try my best to explain :) Good luck!